Test file for the PSTricks 97 distribution: most of the PSTricks examples of The L^{A}TEX Graphics Companion

Sebastian Rahtz — March 1997 Version 1.3 (September 9, 1998)

(a) Example 4-2-1  (b) Example 4-2-2  (c) Example 4-2-3

(d) Example 4-2-4  (e) Example 4-2-5

Figure 1: Section 2

Scary Monsters

(a) Example 4-3-1

(b) Example 4-3-2

Figure 2: Section 3

(a) Example 4-4-1  (b) Example 4-4-2  (c) Example 4-4-3
(d) Example 4-4-4  (e) Example 4-4-5  (f) Example 4-4-6
(g) Example 4-4-7  (h) Example 4-4-8  (i) Example 4-4-9
(j) Example 4-4-10  (k) Example 4-4-11  (l) Example 4-4-12

Figure 3: Section 4 The Buck Stops Here

(a) Example 4-5-1

Beware of The Dog

(b) Example 4-5-2

The dragon ate the women and children

(c) Example 4-5-3

Chapter 1 We go to sea

(d) Example 4-5-4

STOP!

(e) Example 4-5-5

All Hail Caesar!

(f) Example 4-5-6

No Parking

(g) Example 4-5-7

\frac{α}{Ω}

(h) Example 4-5-8

Today’s Menu

(i) Example 4-5-9

If you want to type 3 ∫_{y}^{x} = ∑_{33}^{n=1} and let it follow along a curving line …

(j) Example 4-5-10

Left justified Centred text Right justified

(k) Example 4-5-11

Figure 4: Section 5-A

Tous les souvenirs de naguère_{O mes amis partis en guerre}_{Jaillissent vers le firmament}_{Et vos regards en l’eau dormant}_{Meurent mélancoliquement}_{Où sont-ils Braque et Max Jacob}_{Derain aux yeux gris comme l’aube} ?_{Où sont Raynal Billy Dalize}_{Dont les noms se mélancolisent}_{Comme des pas dans une église}_{Où est Cremnitz qui s’engagea}_{Peut-être sont-ils morts déjà}_{De souvenirs mon âme est pleine}_{Le jet d’eau pleure sur ma peine}_{Ceux qui sont partis à la guerre au nord se battent maintenant}_{Le soir tombeO sanglante mer}_{Jardins où saigne abondamment le laurier rose fleur guerrière}

(a) Example 4-5-13

L^{A}TEX

(b) Example 4-5-14

CHOCOLATE

nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins nuts and raisins

(c) Example 4-5-16

Figure 5: Section 5-B

4

Dog Cat  Mouse

(a) Example 4-6-1

Cat  Dog

(b) Example 4-6-2

Cat  Dog

(c) Example 4-6-3

(d) Example 4-6-4  (e) Example 4-6-5  (f) Example 4-6-6
Cat  Dog

(g) Example 4-6-7

Cat  Dog

(h) Example 4-6-8

Cat  Dog

(i) Example 4-6-9

Cat  Dog

(j) Example 4-6-10 (k) Example 4-6-11

(l) Example 4-6-12

Cat  Dog

(m) Example 4-6-13

Cat  Dog

(n) Example 4-6-14

Cat  Dog

(o) Example 4-6-15

Cat  Dog

(p) Example 4-6-16

Cat  Dog

(q) Example 4-6-17

Cat  Dog

(r) Example 4-6-18

Figure 6: Section 6-A 5

Cat  Dog

(a) Example 4-6-19

Cat  Dog

(b) Example 4-6-20

Cat  Dog

(c) Example 4-6-21

Cat  Dog

(d) Example 4-6-22

Cat

(e) Example 4-6-23

Cat  Dog

(f) Example 4-6-24

Cat  Dog

(g) Example 4-6-25

Figure 7: Section 6-B

1 2 3  1 2 3 4 5  1 2 3 4 5 6 7 8 9 10 11 12

(a) Example 4-6-26

City

Shack  _{House}  Hotel_{a}_{b}_{b}

(b) Example 4-6-27

\begin{matrix}A & B \\ \sqrt{\frac{x+y}{z}} & D \end{matrix}

(c) Example 4-6-28

A  B
\sqrt{\frac{x+y}{z}}  D firstly next then lastly

(d) Example 4-6-29

A  B
\sqrt{\frac{x+y}{z}}  D firstly next then lastly

(e) Example 4-6-30

Waiting Calculating

TRIGGER/activated

[4,11]/finished

(f) Example 4-6-31

Stakeholder  Goal  Criteria _{Sub-goal  Justification Strategy} Model

(g) Example 4-6-32

Figure 8: Section 6-C 7 requirements

design

coding

testing

operations

(a) Example 4-6-33

(b) Example 4-6-34  (c) Example 4-6-35

grandfather uncle cousin father

brother nephew Me son grandson

(d) Example 4-6-36

(e) Example 4-6-37

Figure 9: Section 6-D 8

25  ^{5 L_{a}} a 20 b  15 a  ^{7 L_{baa}} a 5 L_{bab} b_{3 L_{bac}} c_{5 L_{bb}} b

(a) Example 4-6-38

Matilde^{^{Sebastian^{Philip Frederick  Ethel}  ^{Mary Lionel  Agnes}}  ^{Leonor^{Raúl Joaquim  Júlia}  ^{Améliá Alvaro  Augusta}}}

1940 1954

(b) Example 4-6-39

langs german greek h kelly levy doc src

italian russian hyphen

turkish hyphen inputs mf

(c) Example 4-6-40

grandmother aunt cousin

mother sister niece

Me daughter granddaughter

(d) Example 4-6-41

Figure 10: Section 6-E 9

RAMSU

SRawData Scombine ⊕

⊕ SGenerateData

RModes  RNormal  RVertical  RLatched  RTripped  ROther

(a) Example 4-6-42

Monitor Attitude Generator

init

stop

Normal Generator

(b) Example 4-6-43

A→B  B→C  A→D B→E

(c) Example 4-6-44

{A \atop B}

(d) Example 4-6-45

Gabbitas Thring
Gabbitas Thring

(e) Example 4-6-46

Figure 11: Section 6-F

(a) Example 4-7-1

0 1 2 _{0  1  2  3} 0  1  2  3 ; \begin{matrix}0 \\ −1 \\ −2 \end{matrix}  ^{1 {−1 \atop −2} 1 −1}

psaxes(3,2)           psaxes(3,-2)       psaxes(2,1)(0,0)(3,2)

0 ; −1 ; −2 ; −3 0 1 2 0 ; −1 ; −2 ; −3 \begin{matrix}0 \\ −1 \\ −2 \end{matrix}  ^{−1 1 _{0  1}}

psaxes(-3,2)          psaxes(-3,-2)      psaxes(0,0)(-1,0)(1,1)

(b) Example 4-7-2

(()/2)+(()/2)+(()/2)+(()/2)+(()/2)+(()/2)+(()/2)+(()/2)+(()/2)+(()/2)+ French (()/2)+ Latin (()/2)+ Russian (()/2)+ Italian (()/2)+ German (()/2)+ unknown (()/2)+ English

(c) Example 4-7-4

Figure 12: Section 7 11

Words with a shadow Words with a shadow

(a) Example 4-8-1

I Feel Ill!

(b) Example 4-8-2

(c) Example 4-8-3

Viewpoint: -1 -1 1^{FRONT}^{FRONT}^{TOP}  ^{Viewpoint: 1 -1 1^{FRONT}^{BACK}^{TOP}}

Viewpoint: 1 1 1^{BACK}^{BACK}^{TOP}  ^{Viewpoint: -1 1 1^{BACK}^{FRONT}^{TOP}}

(d) Example 4-8-4

Figure 13: Section 8-A

(a) Example 4-8-5

Figure 14: Section 8-B

ππππππππππ

(a) Example 4-9-1

(b) Example 4-9-2

Figure 15: Section 9